APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Tuesday January 11, 2011 (MJD 55573) ---=== OBSERVING TEAM ===--- Howard Brewington Kaike Pan Steph Snedden (afternoon) Support: Paul Harding (Case) ---=== OBSERVING PLAN ===--- MARVELS until 05:37 (4.4 h), then BOSS (7.0 h) ---=== OBSERVING SUMMARY ===--- Tested b side of cartridge 8 with SDSS flat lamps, flux look normal. The first half of the night was partly cloudy. Under the condition, 4 marvels plates were observed. cart 3, plate 4512A. cart 3, plate 4512B. very low signal on fiber 59/60 (fiant object or fiber problem?) cart 1, plate 4475B. cart 8, 4519A. Low signal on fiber 65/66. The 2nd half of the night was plarty to mostly cloudy. 4 BOSS plates were observed. cart 12, plate 4493. cart 16, plate 4509. cart 14, plate 3830. cart 17, plate 3850. We had a netburner problem: 25 min on sky time was lost to the problem (one science exposure failed to read out, plus a few mintues for recovery procedure). ---=== OBSERVING LOG ===--- Afternoon: ---------- 22:50Z - BOSS Collimation checked on cartridge 12. All cameras are in spec. BOSS bias, dark, flat and arc taken. Gcamera dark taken. 23:20Z - Ventilation started. Doors opened. 23:21Z - Cartridge 8 mounted so we could check the blue bundle throughput at the cartridge, however someone offsite was taking sky and other calibration exposures, so we could not move the gang connector. Checkout at the telescope was problem-free. 23:28Z - All actors checked and found to be okay. Night: ------ 00:20Z enclosure off. 00:25Z Last night, MARVELS blue bundle on B side of cartridge 8 were found to have low flux. Cleaning MARVELS gang connectors and did a flux tes. Two exposures (10 sec and 60 sec) were taken, flux levels on all fibers are about the same. In other words, we can not repeat the problem we saw last night. 00:59Z MARVELS calibs T00:59:02.780.fits TIO 349214.00 T01:00:59.760.fits THAR 7317.0771 T01:03:09.550.fits TIO 137264.00 01:00Z cart 3, plate 4512A. Spiral pattern is used to find the field. The final offsets are 5.5" in Ra, -17" in Dec. 01:14Z started 50 science. Partly coudy with 1.3-1.6" seeing. 01:40Z sky imrpoved a little, only some thin clouds. T01:14:19.360.fits STAR 44.682999 Visual data checked good okay. 02:06Z Marvels calibs T02:06:04.970.fits TIO 110251.00 T02:08:01.340.fits THAR 7147.3848 T02:10:11.220.fits TIO 32952.000 02:06Z cart 1, plate 4512B. 02:20 Started science. Cloudy sky, 1.3" seeing. T02:20:20.940.fits STAR 26.977333 Most of time during the exposure, it was heavy cloudy. Signal on fibers 59/60 is very low. 03:12Z Marvels calibs T03:12:06.970.fits TIO 220251.00 T03:14:03.700.fits THAR 7199.2310 T03:16:13.940.fits TIO 247436.00 03:12Z cart 1, plate 4475B. 03:21Z started science. Partly cloudy sky with 1.2-1.4" seeing. T03:21:20.880.fits STAR 45.690334 The data looks good. 04:13Z Marvels calibs T04:13:05.390.fits TIO 195155.00 6013.3154 T04:15:02.150.fits THAR 7308.6152 876.27789 T04:17:12.340.fits TIO 221823.00 04:14Z cart 8, plate 4519A. Sprinal pattern used for the field. The final offsets are: -7" in RA, -15" in Dec. 04:25Z started 50 science. partly cloudy sky. Seeing at 1.7" by the guider. 0.6" by DIMM. T04:25:57.460.fits STAR 35.943668 Low signal on fiber 65/66. 05:17Z MARVELS calibs. T05:17:42.810.fits TIO 167439.00 5949.6997 T05:19:40.580.fits THAR 7072.9229 863.68872 T05:21:50.880.fits TIO 154353.00 05:17Z switch to BOSS. 05:17Z cart 12, plate 4493. 05:33Z started science. Partly cloudy to mostly cloudy sky, seeing at 2.2" by the guider, 0.6" by DIMM. We had a problem: The 3rd exposure on SP2 failed to read out. Reload BOSS controllers recovered it. A following up bias was taken. 6 exposures were taken (but one lost to netburner problem). MJD b1 r1 b2 r2 Texp Sum Plugfile ------------------------------------------------------------ 55573 0.5 1.7 0.4 1.4 900.9 4.0 4493-55573-01 55573 0.6 2.4 0.5 2.2 900.8 5.7 4493-55573-01 55573 0.1 0.2 0.1 0.4 900.8 0.8 4493-55573-01 55573 0.1 0.2 0.2 0.4 900.9 0.9 4493-55573-01 55573 0.4 1.0 0.3 0.6 900.8 2.3 4493-55573-01 ------------------------------------------------------------ 1.7 5.5 1.5 5.0 4504.2 --> 13.7 07:30Z cart 16, plate 4509. 07:37:00 hartSp1: r=31, b=0.1, txt=OK; sp1Temp = 0.2 07:37:23 hartSp2: r=-1338, b=-4.6, txt=OK; sp2Temp = -0.2 07:44Z started science. Sky was mostly cloudy, seeing at 2.0" by the guider. Two exopsures was taken, the plate is not done yet. 08:18Z cart 14, plate 3830. 08:24:49 hartSp1: r=428, b=1.5, txt=OK; sp1Temp = -0.1 08:24:59 hartSp2: r=-477, b=-1.6, txt=OK; sp2Temp = -0.4 08:35Z started science. Mostly cloudy sky, seeing at 2.2" by the guider. 9 exposures were taken. The plate is not done yet. MJD b1 r1 b2 r2 Texp Sum Plugfile ------------------------------------------------------------ 55573 0.4 0.8 0.4 0.8 900.8 2.4 3830-55573-01 55573 0.4 0.8 0.4 0.7 900.9 2.3 3830-55573-01 55573 0.6 0.9 0.6 0.8 900.8 2.9 3830-55573-01 55573 0.3 0.5 0.4 0.5 900.8 1.7 3830-55573-01 55573 0.0 0.0 0.0 0.2 900.8 0.2 3830-55573-01 55573 0.3 0.5 0.3 0.5 900.8 1.6 3830-55573-01 55573 0.4 0.7 0.4 0.7 900.8 2.2 3830-55573-01 55573 0.8 1.2 0.9 1.3 900.8 4.2 3830-55573-01 55573 0.6 1.0 0.8 1.1 900.8 3.5 3830-55573-01 ------------------------------------------------------------ 3.8 6.4 4.2 6.6 8107.3 --> 21.0 11:07Z cart 17, plate 3850. 11:11:53 hartSp1: r=-97, b=-0.3, txt=OK; sp1Temp = 0.3 11:12:06 hartSp2: r=-878, b=-3.0, txt=OK; sp2Temp = 0.1 11:20Z started science. Mostly cloudy sky. Seeing at 1.7" by the guider. Six exposures were taken for the plate. MJD b1 r1 b2 r2 Texp Sum Plugfile ------------------------------------------------------------ 55573 0.2 0.2 0.2 0.1 900.8 0.7 3850-55573-01 55573 0.6 0.8 0.6 0.7 900.8 2.7 3850-55573-01 55573 1.2 1.9 1.3 1.9 900.9 6.3 3850-55573-01 55573 1.1 1.7 1.2 1.7 900.8 5.7 3850-55573-01 55573 0.4 0.6 0.5 0.6 900.8 2.1 3850-55573-01 55573 0.1 0.1 0.1 0.1 900.8 0.4 3850-55573-01 ------------------------------------------------------------ 3.6 5.3 3.9 5.1 5404.9 --> 17.9 13:00Z enclosure on. 13:20Z endNight finished. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): QA Procedures Done (y/n): UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 22:59Z 33 12 1 218 (SW) 1730 - ---=== TELESCOPE STATUS ===--- 23:20Z - Began ventilation. Doors and louvers opened, fans on. 00:20Z - Enclosure off. 13:00Z - Enclosure on, doors and louvers closed, fans off. No interlocks bypassed. Status at:13:20 Z Telescope stowed at: 30,121,00 mount Instrument mounted: Cartridge 16 Counterweights at: 290 lbs LN2 autofill systems: Connected and turned on. 180L LN2 dewar scales: SP1 175 lbs, 18.5 psi SP2 177 lbs, 19 psi ---=== SOFTWARE USED ===--- ----STUI software: STUI: 1.0b13 2010-12-23; alerts: trunk+svn115546; apo: v1_5; boss: v2_3_9; boss.daq: 1.2.14; boss.specMech: 3.0.1; gcamera: v1_2_3; guider: v1_2_9_4; hub: Tag: sdss3_v1_1_5; mcp: v6_6_6; mcp.plc: 26; platedb: v1_0_5; sop: v1_1_5; sos: v1_1_5; sos.idlspec2d: v5_4_24; tcc: 2.14.0; -----Other software sdssProcedure(sdss3): v3_06; idlspec2d (sos3) --- v5_4_24 plate-mapper3 ------ NOCVS:v5_0_7 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -5.5170 -5.5180 Axial B 16.0890 16.0870 Axial C 15.9200 15.9180 Trans D 10.9880 10.9940 Lateral E 10.6807 10.6680 Lateral F 12.2682 12.2682 GALILS Commanded: 10600. -250. 5800. 10250. 13050. 12650. Actual: 10600. -250. 5800. 10250. 13053. 12647. SETMIR VALUES PriDesOrient: 0.00 -13.67 10.53 574.80 742.37 PriOrient: 0.00 -13.67 10.53 574.80 742.37 SECONDARY: ---------- Focus: 0.00 Air Temp.: 0.8 Alt.: 30.153814 MIGS TONIGHT NOMINAL Axial A -7.8350 -7.8360 Axial B -0.1220 -0.1270 Axial C 13.8170 13.8150 Trans D 14.1860 14.1900 GALILS Commanded: 1585381. 1566629. 1598906. -1450. -50. Actual: 1585350. 1566695. 1598880. -1462. -69. SETMIR VALUES SecDesOrient: 1257.00 -9.00 -4.00 -58.00 23.67 SecOrient: 1257.00 -9.02 -4.03 -58.10 23.33 ---=== PROBLEMS IN DETAIL ===---